OpenStudy (anonymous):

If f(x) = -6x - 3, then f(2) = _______?

OpenStudy (anonymous):

-15

OpenStudy (hoblos):

put 2 in the place of x... f(2) = -6(20-3 = -12-3=-15

OpenStudy (anonymous):

put 2 at the place of x